Operations Report – February 9th

Crew 240 Operations Report 09-02-2022

SOL: 9

Name of person filing report: François Vinet

Non-nominal systems: NA

Notes on non-nominal systems: NA

ROVERS

Spirit rover used: no
Hours: 154.8
Beginning charge: NA
Ending charge: NA
Currently charging: managed by Atila

Opportunity rover used: yes
Hours: 83.6
Beginning charge: 100%
Ending charge: 97%
Currently charging: left uncharged, managed by Atila

Curiosity rover used: no
Hours: 166.5
Beginning charge: NA
Ending charge: NA
Currently charging: managed by Atila

Perseverance rover used: yes
Hours: 226.9
Beginning charge: 100%
Ending charge: 95%
Currently charging: left uncharged, managed by Atila

General notes and comments: NA

Summary of Hab operations:

WATER USE: 59 gallons
Water (static tank): 185 gallons
Water (loft tank): 29 gallons
Water Meter: 0155656.2 units
Static to Loft Pump used – yes
Static tank pipe heater (on or off): on
Static tank heater (On or off): on

Toilet tank emptied: yes (we followed the procedure given by Shannon to see if there is some solid matter in the corners of the tank. There is indeed probably some matter in the corners (Clément sent an email with the results). We put a lot of water in the tank, while it was emptying but it is unfortunately not sufficient to fix that problem).

Summary of internet: NA

Summary of suits and radios: To continue about the problem of yesterday about the charge of the batteries, our engineer is still doing some tests about it but the batteries seem ok (at least 3 of them which have been tested).

Summary of GreenHab operations:

WATER USE: 12 gallons at 7:50 am and 8 gallons at 7 pm
Heater: On
Supplemental light: On
Harvest: 290 grams cherry tomatoes

Summary of ScienceDome operations:

Dual split: off
SOC: 53% at 7.00pm

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: A sanitary problem occurred but was fixed with Atila and Shannon.

Questions, concerns and requests to Mission Support: NA